A body is thrown vertically upward from the ground with speed 15m/s and it reaches at the heighest position at time t=1s. Assuming constant air drag, time taken by it to reach the ground from the highest point is nearly (g=10m/s²)
  1. 1s
  2. 1.4s
  3. 0.7s
  4. 1.7s
I am getting my answer as 1.7. Please check if i am correct.

Grade:11

1 Answers

Rishit Dagli
5 years ago
u=15
Let x be the drag
a=-10-x
t=1
v=0
a=v-u/t
10+x = 15
x=5
Now  calculate S
S=ut+0.5at2
S=15+0.5×15
S=22.5 m
During downward journey 
a=15
S=22.5
u=0
t=?
S=ut+0.5at2
22.5=0.5 × 15 × t2
t2=22.5/7.5
t2=3
t=300.5,-30.5
Therfore t = 1.732
Rounding up 
T=1.7 sec
 
Hope it helps 
Rishit  Dagli